Tohban Report 2011-08-31

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 24 August 2011
End Date: 31 August 2011
Tohban: Hazel Bain
Tohban email:
Next Tohban: Lindsay Glesener
List all reports


Solar Activity

Solar activity was low this week with only low C class flares. There are currently 10 active regions on disk, 4 of which will rotate off the disk in the next couple of days. STEREO B shows 1 additional active region due to rotate on disk within the next few days.

How many GOES flares occurred?

 Flares above B, C, M, X class were     29    10     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were     15     8     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were     11     7     0     0

There were RHESSI flares/GOES flares 133 / 39 over the time range 24-Aug-11 31-Aug-11

Memory Management

The SSR is currently around 15%. At the start of the week eclipse data was taken throughout the entire night period but was then reduced to +/- 4 minutes at the start and end of eclipse in preparation for a shutdown at Wallops due to Hurricane Irene. Over the weekend additional passes from Walheim were added to account for the closure at Wallops. Wallops is now back up and running with no damage. A few passes were lost from BGS on the 30th due to technical issues, this has now been fixed. SSR should begin to empty out in the next couple of days.

Data Gaps

There were no significant data gaps this week.

Spacecraft Status

Detector 8 continued to show noise even after the HV was dropped by 6 steps at the start of the week (24th August). The HV was further lowered by 15 steps to 3382V at 23:45 UT 08/28/2011. No further problems have been seen for detector 8.

However detector 4 showed similar noise at 22:30 08/28/2011 followed by a large arc across all detectors. No further noise has been observed and no action has been taken.


Personal tools